OpenStudy (anonymous):

Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

\[f(x)=\frac{ x+a }{ b }\]\[g(x)=cx-d\]

OpenStudy (anonymous):

@waterineyes can you help

OpenStudy (anonymous):

@iGreen ?

OpenStudy (mathmath333):

put any two numbers for a and b say a=1, b=2 \(\large\tt \color{black}{f(x)=\dfrac{x+1}{2}}\) for finding inverse \(\large\tt \color{black}{y=\dfrac{x+1}{2}}\) \(\large\tt \color{black}{2y=x+1}\) \(\large\tt \color{black}{2y-1=x}\) swap x and y \(\large\tt \color{black}{2x-1=y}\) \(\large\tt \color{black}{y=2x-1}\) \(\large\tt \color{black}{g(x)=2x-1}\)